SpaceRef

SpaceRef


ISS On-Orbit Status 13 Feb 2002

Status Report From: NASA HQ
Posted: Wednesday, February 13, 2002

All ISS systems continue to function nominally, except as noted previously or below.

The crew continued preparations for the EVA dry-run on Friday (2/15) and the actual Joint Airlock (JAL) EVA on 2/20. Recharge of two of the new increased-capacity EMU batteries, initiated yesterday, was terminated and the batteries installed in the suit backpacks. Checkout of the EMUs then proceeded on battery power, and it included the liquid-cooled undergarments. The batteries will be topped off tomorrow with the middeck battery charger and then used on the dry-run and EVA. One additional battery was charged today as backup. CDR Onufrienko assisted briefly during the EMU comm check in the beginning. Dan Bursch later also completed an on-board training (OBT) session with the EMU C&W (caution and warning) system and checked out the PGT (pistol grip tool).

FE-1 Carl Walz finished cleaning up after yesterday's change-out of the failed Lab RPCM (remote power controller module) by first installing the protective HDPE (high-density polyethylene) radiation "bricks" around the TeSS (temporary sleep station) at its location at LA1S6 (starboard side closest to aft end cone), and then replacing the TeSS itself including the shear panel, both removed yesterday for the RPCM R&R. [RPCMs are electronically controlled power switches, containing one or more output channels. They are controlled by an internal microcontroller based on EEPROM (electrically erasable programmable read-only memory) microchips. Each output channel is an automatic autonomous circuit breaker (fuse) designed to protect the power switch and downstream wiring. There are six types of RPCM with different output and trip characteristics. They are mounted on coldplates for thermal control and can relatively easily be R&R‚d, once made accessible].

The RPCM failure on 2/11 caused power-down of several Lab systems, including one of the two APS (automated payload switches) and OCA communications, until crew and POC (Payload Operations Center/Huntsville) quickly configured to the other unit (APS-2). Walz today reconfigured the OCA/HDRL (high data rate link) router back to APS-1. The ten-minute task essentially consisted of changing fiber optic cable connections behind UIPs (utility interface panels) on the Lab port side.

CDR Onufrienko ran a test in the SM to evaluate communications flow between the Russian SRK radiation monitoring system and its associated Wiener laptop. The procedure included his use of the 460C DSC (digital still camera) to capture laptop screens displaying SRK data for report to and review by MCC-Moscow (TsUP).

In the payloads/science area, of major significance was the activation and subsequent initial status monitoring of the ADVASC (Advanced Astroculture) plant growth experiment by Bursch and Walz. ADVASC, located in EXPRESS Rack 4 (ER4), is the first facility used to grow plants on ISS, first during Expedition 2 (Usachev, Voss, Helms). [Based on the success of ADVASC-1 (of 91 seeds launched, about 90% germinated), the no. 2 growth chamber came up on Flight UF-1, and it will return to Earth with its results on Mission 8A/STS-110. The study focuses on the mustard plant (Arabidopsis thaliana), world famous in genetic circles. Most significant is the fact that half of the mustard seeds in growth chamber #2 have been harvested from plants grown during ADVASC-1, thus representing a true second generation of space-grown plant life. In the world of genetic research, the flowering mustard plant is the main "model genetics organism", with a fully known genome sequence (completed in 2000). As such, it is a key to identifying genes and determining their functions for entire classes of similar organisms. Its genome contains 25,498 genes encoding proteins from 11,000 families, akin to the functional diversity of the Drosophila fly. ADVASC automatically delivers fluid and nutrients to the root tray and will operate autonomously for 50 to 55 days. The crew will sample nutrients, gases and plant transpiration inside the chamber three times, replenishing the nutrient once at the halfway point, then preserve the samples chemically and store them in the BTR (biotechnology refrigerator)]

. FE-2 Bursch took several pre-EVA readings for the new Canadian EVARM (EVA radiation monitoring) study being conducted by Expedition 4, also for resource tracking (i.e., tracking of power, thermal and hardware resources needed for an EVARM/UOP configuration). During the spacewalk, Bursch and Walz will wear three active EVARM dosimeter badges in pockets sewn into their spacesuits to determine the levels of radiation received to the skin, eyes and blood-forming organs.

CDR Onufrienko performed microbial air sampling using the SZE-MO-21 "Ecosfera" equipment (see yesterday's status).

At 11:26 am EST, the crew participated in a live interactive TV event with school students from the Houston area and Region IV educational service area assembled at another educational conference at the Space Center Houston (SCH), via Ku-band (video) and S-band (audio).

Yuri  and Dan completed the weekly NTXN Interactions data collection, filling out the standard encrypted questionnaire on the MEC (medical equipment computer).

Besides the daily routine maintenance tasks (Lab payload status checks, life support system replacements/inspection, and IMS delta file preparation for downlink), all three crew performed their daily physical exercise on TVIS (treadmill), RED (expander) and VELO (cycle ergometer).

MCC-Houston, aided by U.S. Space Command, is working a predicted conjunction tomorrow, 2/14, with a currently unidentified piece of orbital debris, Object 82120. The encounter will be at 10:24 am EST, at a projected passage distance (range) of 0.82 km (0.5 s.mi.).

Down memory lane:  On 2/20, NASA observes the 40th anniversary of the first U.S. human orbital flight: on that day in 1962 John Glenn became the first American in orbit, circling Earth three times in the one-seater Mercury spacecraft "Friendship 7". And on 2/25, Russia's Space Agency Rosaviakosmos celebrates its 10th anniversary. It's safe to predict that the ISS crew will in some fashion participate in both events.

Today's CEO (crew earth observation) targets were Congo-Zimbabwe Biomass Burning (hot, dry weather has persisted for over a week now, and crew was advised to look to the left of track for smoke plumes and evidence fires being started), European Smog (conditions should have been nearly ideal this pass for heavy smog over the Italian Peninsula and adjacent coastal waters. Of interest during approach to Sicily from the SW: oblique and limb shots to left of track showing the extent of the smog there), E. Mediterranean Dust and Smog (as ISS approached the Balkan Peninsula from the SW, crew was to look well to the right of track over the Aegean Sea for smog plume formations), W. Mediterranean Dust and Smog (with high pressure over the Mediterranean Basin, this may probably was an excellent pass for detecting aerosol plumes over water. Of interest during approach to the Pyrenees Mountains from the SW: aerosol plumes originating over southwestern Europe to the right of track over the Mediterranean Sea), Eastern United States (weather and illumination should be nearly ideal for detecting aerosol accumulations in the atmosphere over the eastern seaboard of the US. As ISS moved northeastward, crew was to shoot oblique and limb views over the coastal waters for best contrast), Gulf of St. Lawrence (it has been some time since the ground has been able to request photos of this target area, and ice season for the Gulf of St Lawrence is now well underway. Light is fading fast this pass, crew was advised to try for near-nadir mapping of the southern gulf near Prince Edward Island), and Tuamotu-Austral Islands (of interest: reefs and atolls between breaks in the clouds during this first pass, using the ESC to map details of what can be seen of the extreme western end of this island chain. Second pass continued northeastward over the western tip of the main Tuamotu Archipelago. Again more clouds than usual but crew was advised to use the ESC to map small reef features where visible.)

U.S. and Russian Segment Status (as of 2:00 pm EST today):

Environmental Control and Life Support (ECLSS):

Elektron O2 generator is On (16 Amp mode). Vozdukh CO2 scrubber is in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is Off. BMP Harmful Impurities unit: Both absorbent beds (Filters #1 & #2) in Purify mode.

  • SM Working Compartment:  Pressure (mmHg) -- 746, temperature (deg C) -- 26.0, ppO2 (mmHg) -- 158.1, ppCO2 (mmHg) -- 2.5 (? see Note below)
  • SM Transfer Compartment:  Pressure (mmHg) -- 747, temperature (deg C) -- 21.0; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • FGB Cabin:  Pressure (mmHg) -- 744, temperature (deg C) -- 22.0; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Node:  Pressure (mmHg) -- 737.53, temperature (deg C) -- 21.8 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab:  Pressure (mmHg) -- 740.17, temperature (deg C) -- n/a, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a;
  • Joint Airlock:  Pressure (mmHg) --  739.87, temperature (deg C) -- n/a; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]. Russian GA (gas analyzer) readings of ppCO2 in the SM are invalid). Electrical Power Systems (EPS):
  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Directed position (125 degrees).
  • SM battery #1 is cycling; all other ( 7) SM batteries are in "partial charge" mode.
  • FGB battery #5 is offline; all other (5) batteries are in "partial charge" mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.
Thermal Control Systems:
  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.
Command & Data Handling Systems:
  • C&C-2 MDM is prime, C&C-3 is back-up, and C&C-1 is in standby.
  • GNC-2 MDM is prime; GNC-1 is Off.
  • LA1 MDM is Off; LA2 and LA3 MDMs are operating.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational. 0—RQúLI>SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.
Communications Systems:
  • S-band is operating nominally, and Ku-band is operating nominally in open loop pointing mode.
  • Audio subsystem operating nominally (except for ATU-2 [audio terminal unit #2]).
  • Video subsystem operating nominally (but VTR2 is non-functional).
  • MCOR (medium-rate communications outage recorder) operating nominally.
Robotics:
  • SSRMS/Canadarm2 at Port stow position (on Keep Alive power on both strings).
  • RWS (robotics workstations) are Off.
ISS Orbit (as of this morning, 7:38 am EST):
  • Mean altitude -- 384.1 km
  • Apogee -- 387.9 km
  • Perigee -- 380.3 km
  • Period -- 92.2 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.0005652
  • Orbits per 24-hr. day -- 15.61
  • Decay rate -- 450 m (mean) in last 24 hours
  • Solar Beta Angle:  -13.2 deg (magnitude increasing)
  • Revolutions since FGB/Zarya launch (Nov. ‚98) -- 18473
  • Current Flight Attitude -- LVLH +XVV ZLV (local vertical/local horizontal: +X-axis in velocity vector; Z-axis in local vertical), with TEA (torque equilibrium attitude [pitch: -10 deg, yaw: -7 deg, roll: 0 deg]). Until 2/14.
For more on ISS orbit and naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter